OpenStudy (anonymous):

The sum of the The sum of the measures of two complementary angles exceeds the difference of their measures by 72°. Find the measure of the smaller angle. of two complementary angles exceeds the difference of their measures by 72°. Find the measure of the smaller angle.

OpenStudy (anonymous):

Let x and y be the angles.Let x>y x+y=90 .......(1) x-y+72=90 x-y=90-72=18 .....(2) adding (1) and (2) and get the value of x and then y

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

yw